OpenStudy (anonymous):

Cart Mass =200kg String compressed by = 5m Initial velocity =25 m/s Find the potential energy stored by the spring while it was compressed.

OpenStudy (anonymous):

So do you know the formula for PE?

OpenStudy (anonymous):

m*g*h

OpenStudy (anonymous):

If the Floor is frictionless then the Kinetic Energy of the Cart will be Converted to P.E of the Spring \[K.E=\frac{ 1 }{ 2 }mv^2\] \[\frac{ 200*25 }{ 2 } = 2500J\]

OpenStudy (anonymous):

So that would be the potential energy?